Barlow Company manufactures three products-A, B, and C. The selling price, variable costs, and contribution margin for one unRequired 1 Required 2 Required 3 Required 4 Calculate the contribution margin per pound of the constraining resource for eachRequired 1 Required 2 Required 3 Required 4 Assuming that Barlow has unlimited demand for each of its three products, what isRequired 1 Required 2 Required 3 Required 4 Assuming that Barlows estimated customer demand is 600 units per product line, wRequired 1 Required 2 Required 3 Required 4 A foreign supplier could furnish Barlow with additional stocks of the raw materia
Answer #1

1)

A B C
Contribution margin per unit [A] 36 72 36
Pounds of direct material required per product [Direct material cost /material cost per pound] [B] 18/9= 2 72/9 = 8 27/9= 3
contribution per pound [A/B] 36/2 =$ 18 Per pound 72/8 =$ 9 per pound 36/3 =$ 12 per pound
Ranking 1 3 2

2)

Since Product A yields highest contribution per constraint resource ,therefore it should be produced to maximum.

Product A
Available Pounds 5400
Number of units of product A that can be produced 5400 pounds /2 pounds per unit = 2700 units

Maximum contribution margin = Number of units of product A that can be produced * contribution per unit

                      = 2700 * 36

                     = $ 97200

3)

Product in sequence of ranking Pounds used to produce maximum quantity(number of units * pounds requirement per unit) Available pounds Contribution margin
A 600 * 2 = 1200 5400-1200= 4200 600*36= 21600
C 600*3=1800 4200-1800= 2400 600*36=21600
B 2400/8=300 units 0 300*72= 21600
Total contribution margin 64800

4)

Since the production requirement of B is not completely met ,The maximum price to pay per pound is equals to contribution margin per pound from product B =$ 9 per pound

> The 9$ per pound in question 4 is not correct. It should be 18$ per pound. The original cost of the material is $9 per pound, the max price it can pay for more materials is 9+9= 18

Lakra Mon, Apr 26, 2021 9:16 PM